OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. One side of a rectangle is 8 inches and the other side is x inches. What values of x will make the perimeter at least 46 inches?

OpenStudy (anonymous):

P=8+8+x+x=16+2x 16+2x>=46

OpenStudy (anonymous):

P=8+8+x+x=16+2x 16+2x>=46 2x>=30 x>=15